when 3 inches are added to each side of a square, its area increases by 57 square inches. What was the perimeter of the initial square?
Answer by ewatrrr(24785)

 
Hi,
Let x represent length of each side of original square
(x+3)^2 = x^2 + 57in^2
x^2 + 6x + 9 = x^2 + 57
6x = 48
x = 8in and P = 4*8in = 32in